OpenStudy (anonymous):

Solve. Give your answer as a mixed numeral. How many pieces of yarn 7 3/4 inches long can be cut from a 48 inch piece of yarn?

OpenStudy (anonymous):

i got 31/192 is this right

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

48 divided by 7 & 3/4 = 48 / (31/4) = 192/31

OpenStudy (anonymous):

my teacher did this 31/4 x 1/48

OpenStudy (anonymous):

he did it wrong ^_^ how many 2 inch pieces of string can be made out of a 10 inch piece? 10/2 = 5

OpenStudy (anonymous):

i think you're right